# Simple Estimation of Semiparametric Models with Measurement Errors

### v3

@@ 3.3 Empirical Illustration: Choice of Transportation Mode
In this section, we illustrate the finite sample properties of the MERM estimator in the context of a classical multinomial choice application: choice of transportation mode (e.g., McFadden, 1974).
@@
To generate the simulated samples, we randomly draw covariates from their joint empirical distribution. To generate the simulated outcomes, we draw $\epsilon_{ij}$ from the standard type-I extreme value distribution. The true value of $\theta_{0}$ is set to be the MLE estimate based on the original dataset. More details about this numerical experiment are given in Appendix H.
@@ 4.1 Multiple Mismeasured Variables
4.2 Weakly Classical and Non-Classical Measurement Errors
In some applications, assuming that the measurement error is classical, i.e., $\varepsilon_{i}$ is independent from $(X_{i}^{*},S_{i})$, may be too restrictive. In this section, we demonstrate how the MERM framework can be used to address this issue. To address general non-classical measurement errors it is convenient to proceed in two steps.
First, in Section 4.2.1, we consider estimation of models with weakly classical measurement errors satisfying $\mathbb{E}\left[\varepsilon_{i}|X_{i}^{\ast},S_{i}\right]=0$. This restriction on $\varepsilon_{i}$ is much weaker than the classical measurement error assumption. For example, the variance of the measurement error $\mathbb{E}\left[\varepsilon_{i}^{2}|X_{i}^{\ast},S_{i}\right]$ (and other moments) could depend on $X_{i}^{\ast}$ and $S_{i}$.
Second, in Section 4.2.2, we consider estimation with general non-classical measurement errors that may be correlated with the mismeasured variable.
4.2.1 Weakly Classical Measurement Errors
In this section, we relax Assumption 2.2 and assume that the measurement error is weakly classical. Specifically, assume that $\mathbb{E}\left[\varepsilon_{i}|X_{i}^{*},S_{i}\right]=0$ or, equivalently, $\mathbb{E}\left[X_{i}|X_{i}^{*},S_{i}\right]=X_{i}^{*}$.
Suppose that $\mathbb{E}\left[\left|\varepsilon_{i}\right|^{k}|X_{i}^{*},S_{i}\right]=O(\tau_{n}^{k})$ for $k\in\{2,\ldots,K\}$. Conditional on $X_{i}^{*}$ and $S_{i}$, the moment function can be expanded as
$\mathbb{E}[g(X_{i},S_{i},\theta)|X_{i}^{\ast},S_{i}]={g(X_{i}^{\ast},S_{i},\theta)}+\sum_{k=2}^{K}\frac{\mathbb{E}\left[\varepsilon_{i}^{k}|X_{i}^{*},S_{i}\right]}{k!}{g_{x}^{(k)}(X_{i}^{*},S_{i},\theta)}+O(\tau_{n}^{K+1}).$ (21)
Suppose $\mathbb{E}\left[\varepsilon_{i}^{k}|X_{i}^{*},S_{i}\right]=v_{k}(X_{i}^{*},S_{i},\omega_{0})$, $k\in\{2,\ldots,K\}$ where $\omega_{0}\in\mathbb{R}^{\dim(\omega)}$ are unknown parameters. Typically, functions $v_{k}$ are assumed to depend only on some of the components of $(X_{i}^{*},S_{i})$. For example, in the nonlinear regression (2), it may be natural for $v_{k}$ to depend on $X_{i}^{*}$ and/or $W_{i}$, but not on $Y_{i}$. Equation (21) then implies that
$\mathbb{E}\left[g(X_{i}^{*},S_{i},\theta)\right]=\mathbb{E}\left[g(X_{i},S_{i},\theta)-\sum_{k=2}^{K}\frac{v_{k}(X_{i}^{*},S_{i},\omega_{0})}{k!}{g_{x}^{(k)}(X_{i}^{*},S_{i},\theta)}\right]+O(\tau_{n}^{K+1}).$ (22)
Equation (22) is an analog of equation (10), which motivates correcting the moment conditions by estimating and subtracting the bias terms $\frac{v_{k}(X_{i}^{*},S_{i},\omega_{0})}{k!}{g_{x}^{(k)}(X_{i}^{*},S_{i},\theta)}$. The corrected moment function takes the form
$\psi(X_{i},S_{i},\theta,\omega)=g(X_{i},S_{i},\theta)-\sum_{k=2}^{K}f_{k}(X_{i},S_{i},\theta,\omega),$ (23)
where
$f_{2}(x,s,\theta,\omega)=$ $\frac{v_{2}(x,s,\omega)}{2}g_{x}^{(2)}(x,s,\theta),\quad f_{3}(x,s,\theta,\omega)=\frac{v_{3}(x,s,\omega)}{6}g_{x}^{(3)}(x,s,\theta),$
$f_{k}(x,s,\theta,\omega)=$ $\frac{v_{k}(x,s,\omega)}{k!}g_{x}^{(k)}(x,s,\theta)-\sum_{\ell=2}^{k-2}\frac{v_{k-\ell}(x,s,\omega)}{(k-\ell)!}\frac{\partial^{k-\ell}}{\partial x^{k-\ell}}f_{\ell}(x,s,\theta,\omega),\text{ for $k\geq 4$}.$
Notice that for $K\leq 3$ the corrected moment functions are similar to the classical EIV case except for $v_{2}(x,s,\omega)$ and $v_{3}(x,s,\omega)$ taking the places of $\mathbb{E}\left[\varepsilon_{i}^{2}\right]$ and $\mathbb{E}\left[\varepsilon_{i}^{3}\right]$ respectively. For $K\geq 4$, we need to bias correct the bias correction terms such as $f_{2}(x,s,\theta,\omega)$ accounting for $v_{2}(X_{i},S_{i},\theta,\omega)$ being evaluated at $X_{i}$ instead of $X_{i}^{*}$. Also, when the conditional moments of $\varepsilon_{i}$ do not depend on $X_{i}^{*}$, the corrected moment function matches the classical EIV case with $v_{k}(s,\omega)$ replacing $\mathbb{E}\left[\varepsilon_{i}^{k}\right]$.
Under the regularity conditions
$\mathbb{E}\left[\psi(X_{i},S_{i},\theta_{0},\omega_{0})\right]=\mathbb{E}\left[g(X_{i}^{*},S_{i},\theta_{0})\right]+O(\tau_{n}^{K+1})=o(n^{-1/2}),$
provided that $O(\tau_{n}^{K+1})=o(n^{-1/2})$.
Thus, the model can be estimated using GMM estimator with the corrected moment function (23), where parameters $\theta_{0}$ and $\omega_{0}$ are estimated together, i.e., the estimator in equation (13) with $\beta\equiv(\theta^{\prime},\omega^{\prime})^{\prime}$. The following example provides a convenient parameterization of functions $v$.
Example 4.1 (Exponential Specification, $K=4$).
Suppose $\varepsilon_{i}=\exp(\omega_{01}X_{i}^{*})\zeta_{i}$, where $\zeta_{i}$ is mean zero and independent of $(X_{i}^{*},S_{i})$. In this case, $\mathbb{E}\left[\varepsilon_{i}^{k}|X_{i}^{*},S\right]=\mathbb{E}\left[\zeta_{i}^{k}\right]\exp(k\omega_{01}X_{i}^{*})$. Then, we can take $v_{k}(x,s,\omega)=\omega_{k}\exp(k\omega_{1}x)$, $k\in\{2,\ldots,4\}$, with $\omega_{0}=(\omega_{01},\mathbb{E}\left[\zeta_{i}^{2}\right],\mathbb{E}\left[\zeta_{i}^{3}\right],\mathbb{E}\left[\zeta_{i}^{4}\right])^{\prime}$. Functions $f_{k}$ in equation (23) are
$f_{2}(x,s,\theta,\omega)$ $=\frac{\omega_{2}}{2}e^{2\omega_{1}x}g_{x}^{(2)}(x,s,\theta),\quad f_{3}(x,s,\theta,\omega)=\frac{\omega_{3}}{6}e^{3\omega_{1}x}g_{x}^{(3)}(x,s,\theta),$
$f_{4}(x,s,\theta,\omega)$ $=\textstyle e^{4\omega_{1}x}\left(\frac{\omega_{4}-6\omega_{2}^{2}}{24}g_{x}^{(4)}(x,s,\theta)-\omega_{1}\omega_{2}^{2}g_{x}^{(3)}(x,s,\theta)-\omega_{1}^{2}\omega_{2}^{2}g_{x}^{(2)}(x,s,\theta)\right).\ \ \hfill\blacksquare$
Identifiability of parameters $\omega_{0}$ depends on the specific model and moment conditions $g$. For the nonlinear regression model with weakly classical EIV Evdokimov and Zeleneev (2022) establish nonparametric identification using instrumental variables.
4.2.2 Non-Classical Measurement Errors
Since $X_{i}^{\ast}$ is not observed, in the absence of some a priori information about its distribution, some assumptions such as $\mathbb{E}\left[\left.X_{i}\right|X_{i}^{\ast},S_{i}\right]=X_{i}^{\ast}$ appear to be necessary, and are routinely imposed in the EIV literature. For example, see Hu and Schennach (2008), Hu and Sasaki (2015), Freyberger (2021), and Schennach (2021) for recent discussions of such assumptions. In the absence of such assumptions one cannot determine even the scale of the true unobserved covariate.
To be specific, suppose the true covariate is denoted by $\mathcal{X}_{i}^{\ast}$ and is mismeasured according to the following model
$X_{i}=\alpha_{0}+\alpha_{1}\mathcal{X}_{i}^{\ast}+\mathcal{E}_{i},\qquad\mathbb{E}\left[\left.\mathcal{E}_{i}\right|\mathcal{X}_{i}^{\ast},S_{i}\right]=0,$ (24)
so $\mathbb{E}\left[\left.X_{i}\right|\mathcal{X}_{i}^{\ast}\right]=\alpha_{0}+\alpha_{1}\mathcal{X}_{i}^{\ast}$ for some unknown $\alpha_{0}$ and $\alpha_{1}>0$. Note that
$\mathbb{E}\left[\left.X_{i}-\mathcal{X}_{i}^{\ast}\right|\mathcal{X}_{i}^{\ast}\right]\neq 0,$
and the measurement error is non-classical. For concreteness, consider the nonlinear regression model:
$\mathbb{E}\left[Y_{i}|\mathcal{X}_{i}^{\ast},W_{i}\right]=\rho\left(\theta_{01}+\theta_{0\mathcal{X}}\mathcal{X}_{i}^{\ast}+\theta_{0W}^{\prime}W_{i}\right)$ (25)
for some known function $\rho$.
Since $\mathcal{X}_{i}^{\ast}$ is unobservable, we can write an observationally equivalent model in terms of another unobservable $X_{i}^{\ast}$ defined as
$X_{i}^{\ast}$ $\equiv\alpha_{0}+\alpha_{1}\mathcal{X}_{i}^{\ast},$ (26)
$\mathbb{E}\left[Y_{i}|X_{i}^{\ast},W_{i}\right]$ $=\rho\left(\tilde{\theta}_{01}+\tilde{\theta}_{0X}{\large X}_{i}^{\ast}+\theta_{0W}^{\prime}W_{i}\right),$ (27)
where $\tilde{\theta}_{0X}\equiv\theta_{0\mathcal{X}}/\alpha_{1},\quad\tilde{\theta}_{01}\equiv\theta_{01}-\tilde{\theta}_{0X}\alpha_{0}$. Thus, parameters $\left(\theta_{01},\theta_{0\mathcal{X}}\right)$ and $\left(\alpha_{0},\alpha_{1}\right)$ are not separately identified without additional information about the location and scale of $\mathcal{X}_{i}^{\ast}$. Note that neither instrumental variables nor multiple measurements of the form (24) provide such information.
Equation (24) and the definition (26) of $X_{i}^{\ast}$ imply that condition $\mathbb{E}\left[\left.X_{i}\right|X_{i}^{\ast},S_{i}\right]=X_{i}^{\ast}$ holds for the observationally equivalent model (27). This observationally equivalent model can be identified and estimated using the approach of Section 4.2.1. This then allows identification of many parameters of interest in the original model (25) because they do not depend on the (unidentified) scale of $\mathcal{X}_{i}^{\ast}$.
First, parameters $\theta_{0W}$ are the same in both models. Moreover, we have $\left(\mathcal{X}_{i}^{\ast}-\mu_{\mathcal{X}^{\ast}}\right)/\sigma_{\mathcal{X}^{\ast}}=\left(X_{i}^{\ast}-\mu_{X^{\ast}}\right)/\sigma_{X^{\ast}}$, and hence the following average structural function
$m(t,w)\equiv\mathbb{E}\left[Y_{i}|\mathcal{X}_{i}^{\ast}=\mu_{\mathcal{X}^{\ast}}+t\sigma_{\mathcal{X}^{\ast}},W_{i}=w\right]$
is identified, since $m(t,w)=\mathbb{E}\left[Y_{i}|X_{i}^{\ast}=\mu_{X^{\ast}}+t\sigma_{X^{\ast}},W_{i}=w\right]$ is identified. As a result, we can identify marginal effects such as $\frac{\partial}{\partial w}\mathbb{E}\left[Y_{i}|\mathcal{X}_{i}^{\ast}=\mu_{\mathcal{X}^{\ast}},W_{i}=w\right]=\frac{\partial}{\partial w}m(0,w)$, or the effect of increasing $\mathcal{X}_{i}^{\ast}$ from $\mu_{\mathcal{X}^{\ast}}$ to $\mu_{\mathcal{X}^{\ast}}+\lambda\sigma_{\mathcal{X}^{\ast}}$, i.e., $m(\lambda,w)-m(0,w)$. The corresponding marginal effects averaged with respect to the distribution of $\left(\mathcal{X}_{i}^{\ast},W_{i}\right)$ are also identified and can be estimated following Remark 3. Thus, in this model, the MERM estimator from Section 4.2.1 allows estimation of many economically meaningful marginal effects, even though the true measurement model is given by equation (24), and the measurement error $X_{i}-\mathcal{X}_{i}^{\ast}$ is non-classical. [^10] In the constructed observationally equivalent model (27) we have $\varepsilon_{i}=\mathcal{E}_{i}$, and the MERM approach is justified when $\sigma_{\mathcal{E}}^{2}/\sigma_{X}^{2}$ is moderate. The true measurement error $X_{i}-\mathcal{X}_{i}^{\ast}$ does not need to be small.
At the same time, identification of the true values of $\theta_{01}$ and $\theta_{0\mathcal{X}}$ or some marginal effects such as $\frac{\partial}{\partial\varkappa}\mathbb{E}\left[Y_{i}|\mathcal{X}_{i}^{\ast}=\varkappa,W_{i}=w\right]$ requires identifying $\alpha_{0}$ and $\alpha_{1}$.
Using the summary statistics for $\mathcal{X}^{\ast}$
In some applications the researchers may have information about the marginal distribution of $\mathcal{X}^{\ast}$ such as its mean $\mu_{\mathcal{X}^{\ast}}$ and variance $\sigma_{\mathcal{X}^{\ast}}^{2}$, e.g., from an administrative dataset. [^11] It is important that $\mathcal{X}_{j}^{\ast}$ and $\left(X_{i},S_{i}\right)$ are drawn from the same population., [^12] Knowing $\sigma_{\mathcal{X}^{\ast}}^{2}$ does not identify even the unconditional variance $\sigma_{\varepsilon}^{2}$, since $\sigma_{X}^{2}=\alpha_{1}^{2}\sigma_{\mathcal{X}^{\ast}}^{2}+\sigma_{\varepsilon}^{2}$. Then, the MERM approach can be used to estimate $\alpha_{0}$ and $\alpha_{1}$ together with parameters $\theta_{0}$. For any original moment condition $\mathbb{E}\left[g\left(\mathcal{X}_{i}^{\ast},S_{i},\theta_{0}\right)\right]=0$ consider the augmented moment condition $\mathbb{E}\left[g_{A}\left(\mathcal{X}_{i}^{\ast},S_{i},\theta_{0}\right)\right]=0$, where
$g_{A}\left(\mathcal{X}_{i}^{\ast},S_{i},\theta_{0}\right)\equiv\left(g\left(\mathcal{X}_{i}^{\ast},S_{i},\theta_{0}\right)^{\prime},\mathcal{X}_{i}^{\ast}-\mu_{\mathcal{X}^{\ast}},\left(\mathcal{X}_{i}^{\ast}-\mu_{\mathcal{X}^{\ast}}\right)^{2}-\sigma_{\mathcal{X}^{\ast}}^{2}\right)^{\prime}.$
Define $X_{i}^{\ast}\equiv\alpha_{0}+\alpha_{1}\mathcal{X}_{i}^{\ast}$, $\tilde{\theta}_{0}\equiv\left(\theta_{0}^{\prime},\alpha_{0},\alpha_{1}\right)^{\prime}$, and $\tilde{g}_{A}(X_{i}^{\ast},S_{i},\tilde{\theta}_{0})\equiv g_{A}\left(\left.\left(X_{i}^{\ast}-\alpha_{0}\right)\right/\alpha_{1},S_{i},\theta_{0}\right)$. Then one can take $\mathbb{E}[\tilde{g}_{A}(X_{i}^{\ast},S_{i},\tilde{\theta}_{0})]=0$ as the new “original” moment condition (1). Since the condition $\mathbb{E}\left[X_{i}|X_{i}^{\ast},S_{i}\right]=X_{i}^{\ast}$ holds, $\tilde{\theta}_{0}$ can be estimated using the MERM approach and the corrected moments (23) applied to $\tilde{g}_{A}$ in place of $g$.
Evdokimov and Zeleneev (2022) show how an identification argument based on the assumption of weakly classical measurement errors can be used for nonparametric identification in the settings with non-classical measurement errors. The linearity of equation (24) simplifies the estimation procedure, but the underlying approach is valid nonparametrically. Using a parsimonious model, this section provides a practical way of dealing with non-classical measurement errors in empirical applications, avoiding any nonparametric estimation.

### v4

@@ 3.2 Estimation and Inference in a Multinomial Choice Model
To further check the limits of applicability of our method, we also consider larger values of $\tau\in\{1,3/2,2\}$. The numerical results analogous to the ones reported in Table LABEL:tab:_MC_mult_logit_2 are provided in Table 6 in Appendix F.1. Specifically, we find that inference results based on the correction scheme with $K=4$ remain accurate even for $\tau=1$. Unsurprisingly, inference becomes less reliable for bigger $\tau=3/2$ and $\tau=2$. At the same time, while the $K=4$ correction scheme fails to entirely eliminate the EIV bias in these designs, it still removes a big fraction of the bias and greatly improves on MLE in terms of the RMSE. Thus, while inference based on our estimator might be less reliable in extreme settings when the measurement error overwhelms the signal, the MERM estimator using $K=4$ appears to be sufficiently accurate over a wide range of $\tau$.
@@ 3.3 Empirical Illustration: Choice of Transportation Mode
In this section, we illustrate the finite sample properties of the MERM estimator in the context of a classical multinomial choice application: choice of transportation mode (e.g., McFadden1974JPubE).
@@
To generate the simulated samples, we randomly draw covariates from their joint empirical distribution. To generate the simulated outcomes, we draw $\epsilon_{ij}$ from the standard type-I extreme value distribution. The true value of $\theta_{0}$ is set to be the MLE estimate based on the original dataset. More details about this numerical experiment are given in Appendix I.
@@ 4 Extensions
4.1 Data-driven choice of $K$
Making an appropriate choice of the expansion order $K$ is important for the estimation procedure. One has to be cautious not to take $K$ too small, as this may result in an estimator that only partially removes the EIV bias. On the other hand, picking a larger $K$ than needed might inflate standard errors and result in less powerful inference.
In this section, we address this issue by providing a data-dependent procedure for selecting $K$. We demonstrate that our procedure has desirable theoretical properties. We also find that the procedure has good finite sample properties in a set of Monte Carlo simulation experiments across different values of $\tau$ and sample sizes.
Consider two alternative values of the expansion order: $L$ and $K$, where $2\leq L<K$. In practice, even-order biases tend to dominate, so to reduce the set of choices, it is useful to focus on even values of the expansion orders $L$ and $K$. For example, one would typically be interested in choosing between $L=2$ and $K=4$.
Let $\hat{\beta}_{L}$ and $\hat{\beta}_{K}$ denote the corresponding MERM estimators. The estimator $\hat{\beta}_{L}$ should be preferred as having smaller asymptotic variance provided that its remaining EIV bias is negligible relative to its standard error. Otherwise, the more conservative $\hat{\beta}_{K}$ should be used instead.
Note that Lemma 1 suggests that the remaining asymptotic bias of $\hat{\beta}_{L}$ (due to the additional terms accounted for when the expansion of higher order $K$ is used) is given by (up to an $o(n^{-1/2})$ remainder)
$\text{AsyB}(\hat{\beta}_{L})\equiv B\sum_{k=L+1}^{K}\gamma_{0k}\mathbb{E}[g_{x}^{(k)}(X_{i},S_{i},\theta_{0})],\quad B\equiv-(\Psi^{\prime}\Xi\Psi)^{-1}\Psi^{\prime}\Xi,$
where matrix $B$ is based on the moments used for estimation of $\hat{\beta}_{L}$.
Importantly, $\gamma_{0k}=O\left(\sigma_{\varepsilon}^{k}\right)$, and hence for $k>2$ we can estimate a bound on $\sqrt{n}\gamma_{0k}$ sufficiently quickly to provide a valid procedure for choosing $K$. To this end, we first estimate the model using the larger $K$. Let $\hat{\beta}_{K}=(\hat{\theta}^{\prime},\hat{\gamma}^{\prime})^{\prime}$ and $\hat{\sigma}_{\varepsilon}^{2}\equiv 2\hat{\gamma}_{2}$, where we dropped the additional subscripts $K$ for notation simplicity. Then, we can estimate $\sigma_{\varepsilon}^{K}$ by $\hat{\sigma}_{\varepsilon}^{K}\equiv(\hat{\sigma}_{\varepsilon}^{2})^{K/2}$. Using $\hat{\sigma}_{\varepsilon}^{2}=\sigma_{\varepsilon}^{2}+O_{p}(n^{-1/2})$, in the appendix we show that
$\sqrt{n}\left(\hat{\sigma}_{\varepsilon}^{K}-\sigma_{\varepsilon}^{K}\right)=O_{p}\left(\sigma_{\varepsilon}^{K-2}+n^{-\left(K-2\right)/4}\right)=o_{p}\left(1\right).$ (19)
Next, consider a sequence $\varkappa_{n}\rightarrow 0$, which we will specify precisely later, and let
$\delta_{n}=1\left\{\max_{1\leq\ell\leq\dim(\beta)}\left|\hat{\Sigma}_{\ell\ell}^{-1/2}\sqrt{n}\hat{\sigma}_{\varepsilon}^{K}\hat{B}_{\ell\cdot}\overline{g}_{x}^{(K)}(\hat{\theta})\right|\leq c_{K}\varkappa_{n}\right\},$ (20)
where $\hat{\Sigma}$ and $\hat{B}$ are consistent estimators of the asymptotic variance of $\hat{\beta}_{L}$ and of $B$, with $\hat{\Sigma}_{\ell\ell}$ denoting its $\ell$-th diagonal element, $\hat{B}_{\ell\cdot}$ denoting its $\ell$-th row, ${\overline{g}_{x}^{(K)}(\hat{\theta})\equiv n^{-1}\sum_{i=1}^{n}g^{(K)}(X_{i},S_{i},\hat{\theta})}$, and $c_{K}>0$ is a constant that we will calibrate below after we state the main theoretical result of this section.
If $\delta_{n}=1$, the researchers should select $\hat{\beta}_{L}$. Otherwise, $\hat{\beta}_{K}$ should be used. The following lemma demonstrates that the proposed selection procedure has desirable theoretical properties.
Lemma 3.
Suppose the hypotheses of Theorem 2 hold for some even $K\geq 4$, and consider $L$ satisfying $2\leq L<K$. Suppose $\varkappa_{n}n^{\left(K-L-1\right)/\left(2L+2\right)}\rightarrow 0$. Then $\sqrt{n}\text{AsyB}(\hat{\beta}_{L})\delta_{n}=o_{p}\left(1\right)$.
Moreover, consider $\varkappa_{n}=n^{-\left(K-L-1\right)/\left(2L+2\right)}\left(\ln n\right)^{-a}$ for any $a>0$. Then the criterion is consistent, in the sense that if $\tau_{n}=o(n^{-\frac{1}{2L+2}-\epsilon})$ for any $\epsilon>0$, the criterion will choose $\hat{\beta}_{L}$ with probability approaching one.
The first part of Lemma 3 shows that, provided that $\varkappa_{n}$ goes to zero sufficiently fast, $\hat{\beta}_{L}$ is selected (i.e., $\delta_{n}=1$) only when its asymptotic bias is negligible. Next, note that $\hat{\beta}_{L}$ is asymptotically unbiased as long as $\tau_{n}=(n^{-\frac{1}{2L+2}})$. The second part of the lemma shows that, for the suggested choices of $\varkappa_{n}$, the criterion is non-vacuous, i.e., that it does select the MERM estimator with a smaller expansion order $L$ when this is appropriate. Typically, one would pick L=K2L=K-2, so the lemma suggests taking $\varkappa_{n}=n^{-1/\left(2K-2\right)}\left(\ln n\right)^{-a}$ in this case.
In practice, it is important to pick an appropriate constant $c_{K}$ used in the construction of $\delta_{n}$ in equation (20). When constructing $\delta_{n}$, we used $\left|\gamma_{0K}\right|\propto\sigma_{\varepsilon}^{K}$, which motivates choosing $c_{K}=\sigma_{\varepsilon}^{K}/\left|\gamma_{0K}\right|$. It is convenient to use a rule-of-thumb approach, using a reference distribution to determine $c_{K}$. It turns out that the normal distribution is not only convenient, but also sufficiently conservative (note that the bigger $\gamma_{0K}$ is, the smaller $c_{K}$ is, resulting in a more conservative selection procedure). For example, suppose $K=4$, and consider using Student’s $t(\nu)$ distribution as the reference distribution. Then, for $\nu\geq 5$, the biggest $\left|\gamma_{04}\right|/\sigma_{\varepsilon}^{4}$ and the smallest $c_{4}$ correspond to $\nu=\infty$ matching the normal distribution. [^11] Note that Student’s $t(\nu)$ distribution does not have a finite $5$-th moment $\nu\leq 5$.
Since for the normal distribution we have $\gamma_{0K}=\sigma_{\varepsilon}^{K}/K!!$ for even $K$, we recommend using $c_{K}=K!!$ in equation (20). Finally, while we recommend using normal distribution as the reference distribution, we also stress that the results of Lemma 3 hold even if the measurement error is not normal or if it is skewed.
We summarize the proposed procedure in the following suggested algorithm.
Algorithm 1.
  1. 1.
    Pick a large enough even $K\geq 4$ and compute $\hat{\beta}_{K}=(\hat{\theta}^{\prime},\hat{\gamma}^{\prime})^{\prime}$ and $\hat{\sigma}_{\varepsilon}^{2}=2\hat{\gamma}_{2}$.
  2. 2.
    Compute $\delta_{n}$ in (20) with L=K2L=K-2, $c_{K}=K!!$, and $\varkappa_{n}=\left(n\ln n\right)^{-1/\left(2K-2\right)}$.
  3. 3.
    Pick the length of expansion L=K2L=K-2 if $\delta_{n}=1$, otherwise keep the initial $K$.
This procedure can be iterated if desired.
To illustrate the performance of the algorithm provided above, we revisit the numerical experiment considered in Section 3.2. We focus on choosing between $K=4$ and $L=2$, and, as in Section 3.2, we report results for $n=2000$ in Table 5 below. To ensure that the proposed algorithm performs well in a variety of sample sizes, we also consider $n=1000$ and $n=4000$ and report the corresponding results in Tables 7 and 8 in Appendix F.2.
We report the finite sample bias and RMSE for the naive MLE estimator, as well as for the MERM estimators using $K=2$ and $K=4$, and for the adaptive MERM estimator using data-driven $K$ following Algorithm 1. Notice that for the considered sample sizes, $K=2$ is preferred when $\tau=1/4$, and $K=4$ is preferred when $\tau=3/4$. We find that in both of these regimes and for all sample sizes, the adaptive estimator using data-driven $K$ is essentially equivalent to the preferred estimators, i.e., our procedure selects the appropriate $K$. Interestingly, in the intermediate regime with $\tau=1/2$, the adaptive estimator also has the smallest bias among the considered estimators, coming at the cost of a slightly bigger RMSE compared to the MERM estimator using $K=4$. Thus, the considered numerical experiment suggests that our algorithm has good finite sample properties supporting the findings of Lemma 3.
Table 5: Choice of $K$ simulation results for the multinomial logit model, $n=2000$
MLE $K=2$ $K=4$ data-driven $K$
bias, $10^{-2}$ rmse, $10^{-2}$ bias, $10^{-2}$ rmse, $10^{-2}$ bias, $10^{-2}$ rmse, $10^{-2}$ bias, $10^{-2}$ rmse, $10^{-2}$
$\tau=1/4$
$\partial p_{1}/\partial x$ -3.24 3.51 0.74 2.74 1.13 2.95 0.75 2.75
$\partial p_{1}/\partial w_{1}$ 2.32 2.84 -0.11 2.30 -0.31 2.31 -0.11 2.30
$\partial p_{1}/\partial w_{2}$ 0.48 0.90 -0.04 0.87 -0.08 0.87 -0.04 0.87
$\partial p_{2}/\partial x$ 1.96 2.28 -0.40 1.92 -0.63 2.03 -0.41 1.93
$\partial p_{2}/\partial w_{1}$ -1.16 1.42 0.06 1.15 0.15 1.16 0.06 1.15
$\partial p_{2}/\partial w_{2}$ -0.96 1.78 0.09 1.74 0.17 1.75 0.09 1.74
$\partial p_{0}/\partial x$ 1.28 1.63 -0.34 1.47 -0.50 1.56 -0.34 1.48
$\partial p_{0}/\partial w_{1}$ -1.16 1.42 0.06 1.15 0.15 1.16 0.06 1.15
$\partial p_{0}/\partial w_{2}$ 0.48 0.88 -0.05 0.87 -0.09 0.88 -0.05 0.87
$\tau=1/2$
$\partial p_{1}/\partial x$ -8.97 9.04 -1.69 3.10 0.97 3.05 0.88 3.15
$\partial p_{1}/\partial w_{1}$ 6.44 6.62 1.39 2.81 -0.21 2.42 -0.15 2.48
$\partial p_{1}/\partial w_{2}$ 1.28 1.47 0.29 0.94 -0.06 0.92 -0.05 0.93
$\partial p_{2}/\partial x$ 5.22 5.31 1.05 2.16 -0.53 2.15 -0.48 2.20
$\partial p_{2}/\partial w_{1}$ -3.21 3.30 -0.69 1.40 0.10 1.21 0.08 1.24
$\partial p_{2}/\partial w_{2}$ -2.52 2.88 -0.56 1.87 0.13 1.84 0.11 1.85
$\partial p_{0}/\partial x$ 3.75 3.85 0.64 1.59 -0.44 1.65 -0.40 1.68
$\partial p_{0}/\partial w_{1}$ -3.23 3.32 -0.70 1.41 0.10 1.21 0.08 1.24
$\partial p_{0}/\partial w_{2}$ 1.23 1.41 0.28 0.93 -0.07 0.92 -0.06 0.93
$\tau=3/4$
$\partial p_{1}/\partial x$ -13.35 13.38 -6.83 7.32 0.71 3.29 0.71 3.29
$\partial p_{1}/\partial w_{1}$ 9.69 9.80 4.95 5.61 0.01 2.62 0.01 2.62
$\partial p_{1}/\partial w_{2}$ 1.81 1.94 1.01 1.35 -0.01 0.98 -0.01 0.98
$\partial p_{2}/\partial x$ 7.48 7.52 4.06 4.45 -0.37 2.35 -0.37 2.35
$\partial p_{2}/\partial w_{1}$ -4.83 4.88 -2.47 2.81 -0.01 1.31 -0.01 1.31
$\partial p_{2}/\partial w_{2}$ -3.51 3.76 -1.99 2.66 0.03 1.97 0.03 1.97
$\partial p_{0}/\partial x$ 5.87 5.92 2.77 3.14 -0.34 1.80 -0.34 1.80
$\partial p_{0}/\partial w_{1}$ -4.87 4.93 -2.48 2.81 -0.01 1.31 -0.01 1.31
$\partial p_{0}/\partial w_{2}$ 1.70 1.82 0.98 1.31 -0.02 0.99 -0.02 0.99
  • This table reports the simulated finite sample bias and RMSE of the MLE and the MERM estimators for the partial derivatives $\partial p_{j}(x,w,\theta_{0})/\partial x$, $\partial p_{j}(x,w,\theta_{0})/\partial w_{1}$, $\partial p_{j}(x,w,\theta_{0})/\partial w_{2}$ for $j\in\{1,2,0\}$ evaluated at the population mean. The true values of the marginal effects are $(\partial p_{1}/\partial x,\partial p_{2}/\partial x,\partial p_{0}/\partial x)=(0.222,-0.111,-0.111)$ and zeros for the rest. The results are based on 5,000 replications.